# Batch Bayesian Optimization for High-Dimensional Experimental Design: Simulation and Visualization  [^1] A footnote for the title

### v1

@@ 2.3 Benchmark Approach
We benchmark BO with the two test objective functions (A($\bf X$) and H($\bf X$)), two acquisition functions (EI and UCB) with respective hyperparameters, as well as three batch-picking methods (LP, KB, and CL), without or with different levels of noise according to the workflow illustrated in Fig. 2. The workflow involves initial sampling (Fig. 2(a)), repeating the BO iterations which include evaluating the objective values at each sampled point (Fig. 2(b)), model training (Fig. 2(c)), and suggesting the input points to evaluate for the next batch of points (Fig. 2(d)), and evaluating the learning curves after the end of BO, in our case after 50 iterations (Fig. 2(e)). BO settings, apart from the benchmark variables, are kept the same across the benchmark simulations, as detailed next.
@@
$y^{i}=\rm f({\bf X}^{i})+\epsilon^{i},\epsilon^{i}\sim\textit{N}(0,(\sigma^{i})^{2})$ (7)
We introduce noise in two distinct methods to analyze its impact on BO. Following the common practice in the literature 25, noise was added as a percentage of the maximum GT value,
@@
$\rm\sigma^{i}=\textrm{(kernel\; amplitude)}_{\epsilon=0}\times{\rm proportion\;of\;noise}$ (9)
@@
Model training (Fig. 2(c)) involves training a GPR surrogate model at each iteration of BO. In each scenario, the ARDMatern52 is used as the kernel function. Automatic relevance determination (ARD) kernels assume a different length scale for each dimension and tune them separately. 41, 42 In a previous benchmark by Liang et al., the ARD kernels performed better than the non-ARD ones.1 The ARD Matern kernel function is:
@@
where $\sigma^{2}$ is the amplitude parameter, L is a length scale parameter, $\tau(.)$ is the gamma function, $\rm K_{\nu}$ is a modified Bessel function of the second kind and $\nu$=5/2 refers to the smoothness of the function (lower value means more smooth).5, 43 The kernel amplitude and length scales are hyperparameters that must be correctly tuned to construct a realistic GPR surrogate model. Kernel amplitude represents the value range in the surrogate model predictions, and length scale represents the correlation between two points in the input space $\bf X$ of the surrogate model. Additionally, GPR model has a Gaussian noise variance (GNV) hyperparameter, representing the uncertainty associated with each observation and, when applied to experiments, reflecting the noise level in the data. Therefore, the GPR surrogate model may overly smooth the observed data if the noise variance is too high, whereas the model may overfit the noisy observations if the noise variance is too low. Both cases lead to suboptimal decisions. All three hyperparameters are autotuned in the simulation and tracked with iterations.
The next four points to sample are suggested according to the acquisition function and batch-picking method (Fig. 2(d)). We test LP, KB, and CL as batch-picking methods for both Ackley and Hartmann functions with both EI and UCB as the acquisition functions with optimized hyperparameters based on performance metrics. We find LP outperforms KB and CL for both functions independent of the choice of acquisition function. For details see Supporting Information appedix section D. For the rest of the paper, LP is used as the batch-picking method unless explicitly specified.
@@ 2.4 Metrics for Performance Evaluation During Benchmarking
To characterize the optimization of the BO results, we compute instantaneous regret (IR) for ${\bf X}$ and $y$: ${\rm IR}({\bf X}_{k})=||{{\bf X}_{k}^{50}}^{*}-{\bf X}_{\rm max}||$, and ${\rm IR}(y_{k})=|\mu_{\rm D}({{\bf X}_{k}^{50}}^{*})-y_{\rm max}|$, and then average the IR over all 99 LHS samplings to establish statistical variation:
$\langle{\rm IR}({\bf X})\rangle=\frac{\sum_{k=1}^{99}||{{\bf X}_{k}^{50}}^{*}-{\bf X}_{\rm max}||}{99}$ (12)
@@ 3 Results and Discussion
3.1 Noise-free
3.1.1 Ackley
Noise-free optimizations are investigated first to determine the ideal performance of BO on our optimization tasks, and to illustrate the convergence of the BO in terms of our performance evaluation metrics. Fig. 3 shows the BO of the Ackley function using the UCB acquisition policy. Fig. 3(a) shows $||{\bf X}_{k}^{i}-{\bf X}_{\rm max}||$, the deviation of the current optimal location from the GT maximum location (${\bf X}_{\rm max}$) as a function of iteration index $i$ for all 99 LHS samplings. Since the initial pick is random, none of the 99 samplings yield initial vectors close to ${\bf X}_{\rm max}$. Thus, $||{\bf X}^{1}-{\bf X}_{\rm max}||>15$ which is not surprising since the search space is a large hypercube measuring $[-32.768,32.768]^{6}$. After 10 iterations significant progress towards the optimum is made, $||{\bf X}^{10}-{\bf X}_{\rm max}||<10$ for all 99 samplings. At the last iteration ${\rm IR}({\bf X}_{k})=||{\bf X}_{k}^{50}-{\bf X}_{\rm max}||<1$ and the estimated optimal inputs for all 99 samplings are very close to the GT maximum.
@@ 3.1.1 Ackley
font=footnotesize Refer to caption
Figure 3: BO results using UCB with $\beta$ = 1 on noise-free Ackley function. Learning curve in (a) $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for all 99 LHS initial starts. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representation (x1 vs. x2) for the 50th percentile BO model, showing how BO iterations zero in on ${\bf X}_{\rm max}$ (cyan cross). Light blue circles are the 24 initial LHS selections and the pink to dark red points are training points (progressively darker). (d) Zoomed-in 2D heat map for variables x1 and x2 within the range [-2, 2] near ${\bf X}_{\rm max}$. (e) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
@@
Fig. 3(b) shows the $y$ value estimated by the surrogate GPR model, $\mu_{\rm D}({\bf X}^{*})$, as a function of BO iterations. At iteration 1, all $y$-value estimates are below $-15$, whereas around iteration 20 estimates have improved to be above $-5$, and at the final iteration the $y$-values are all close to 0, the maximum $A({\bf X})$ value. Fig. 3(c) shows a 3D representation of the surrogate model and scatter plot of the inputs of the 50th percentile sampling on two of the six ${\bf X}$ dimensions. Along the input range from -32.768 to 32.768, the objective function ranges from - 22.3 to 0. Each dot indicates an input point; blue dots indicate initial LHS points and the red dots from light to dark represents the learning progression. As iterations progress more points are sampled close to ${\bf X}_{\rm max}$, indicating that the BO exploits the optimum region. Fig. 3(d) shows the 2D heat map with focused domain range near maximum $A({\bf X})$ and reveals that most of the points for the Ackley function are sampled around ${\bf X}_{\rm max}$.
In addition to the analysis of performance metrics, the evolution of GPR hyperparameters is provided in Fig. S3. We observe that all length scales converged to similar values as expected from the rotational symmetry around the origin of the Ackley function. It should be noted that the exploration hyperparameter $\beta$ greatly affects BO convergence.34, 12 Table  S1 shows the different metrics, $\rm\langle IR({\bf X})\rangle$, $\rm\langle CR({\bf X})\rangle$, $\rm\langle IR(y)\rangle$, and $\rm\langle CR(y)\rangle$, all normalized to the ${\bf X},y$ ranges, for different $\beta$ values. We chose the $\beta$ that produces the smallest $\rm\langle IR(\bf X)\rangle$ which we determine as $\beta=1$, which also yields the smallest $\langle{\rm CR}({\bf X})\rangle$ and $\langle{\rm CR}(y)\rangle$.
@@
3.1.2 Hartmann
@@
Fig. 5a shows $||{\bf X}_{k}^{i}-{\bf X}_{\rm max}||$ for the Hartmann test function using the UCB acquisition policy. Initially, the distance to the maximum $||{\bf X}^{1}_{k}-{\bf X}_{\rm max}||$ ranges from $0.2$ to $1.5$ in the unit hypercube. The distance to ${\bf X}_{\rm max}$ for the Hartmann function show relatively larger variance compared to the Ackley, which we attribute to those LHS sampling result in finding a optimum close to ${\bf X}_{\rm max,2}$ rather than ${\bf X}_{\rm max}$. In Fig. 5(a), for $\approx 75\%$ of the samplings, $||{\bf X}_{k}-{\bf X}_{\rm max}||$ converges to 0 while for the remaining $\approx 25\%$, $||{\bf X}_{k}-{\bf X}_{\rm max}||$ converges to $||{\bf X}_{\rm max,2}-{\bf X}_{\rm max}||\approx 1.1$. In some LHS samplings ${\bf X}_{k}$ is initially near ${\bf X}_{\rm max,2}$ but still ends up at ${\bf X}_{\rm max}$. Overall, because of the second maximum, it is significantly more challenging to find the GT max in the Hartmann landscape compared to the Ackley landscape.

### v3

@@ 3.3 Benchmark Approach
We benchmark batch BO using Emukit open source BO code package [45, 46] with the two test objective functions (A($\bf X$) and H($\bf X$)), two acquisition functions (EI and UCB) with respective hyperparameters, as well as three analytical batch-picking methods (LP, KB, and CL), with different levels of noise according to the workflow illustrated in Fig. 2. The workflow involves initial sampling (Fig. 2(a)), repeating the BO iterations which include evaluating the objective values at each sampled point (Fig. 2(b)), model training (Fig. 2(c)), and suggesting the input points to evaluate for the next batch of points (Fig. 2(d)), and evaluating the learning curves after the end of BO, in our case after 50 iterations (Fig. 2(e)). BO settings, apart from the benchmark variables, are kept the same across the benchmark simulations, as detailed next.
@@
$y^{i}=\rm f({\bf X}^{\textit{i}})+\epsilon^{\textit{i}},\epsilon^{\textit{i}}\sim\textit{N}(0,(\sigma^{i})^{2})$ (7)
We introduce noise in two distinct methods to analyze its impact on BO. Following the common practice in the literature [8], noise was added as a percentage of the maximum GT value,
@@
$\sigma^{i}=\rm\textrm{(kernel\; amplitude)}_{\epsilon=0}\times{\rm proportion\;of\;noise}$ (9)
@@
Model training (Fig. 2(c)) involves training a GPR surrogate model at each iteration of BO. In each scenario, the ARDMatern52 is used as the kernel function. Automatic relevance determination (ARD) kernels assume a different length scale for each dimension and tune them separately. [47, 48] In a previous benchmark by Liang et al., the ARD kernels performed better than the non-ARD ones.[4] The ARD Matern kernel function is:
@@
where $\sigma^{2}$ is the amplitude parameter, L is a length scale parameter, $\tau(.)$ is the gamma function, $\rm K_{\nu}$ is a modified Bessel function of the second kind and $\nu$=5/2 refers to the smoothness of the function (lower value means more smooth).[3, 49] The kernel amplitude and length scales are hyperparameters that must be correctly tuned to construct a realistic GPR surrogate model. Kernel amplitude represents the value range in the surrogate model predictions, and length scale represents the correlation between two points in the input space $\bf X$ of the surrogate model. Additionally, GPR model has a Gaussian noise variance (GNV) hyperparameter, representing the uncertainty associated with each observation and, when applied to experiments, reflecting the noise level in the data. Therefore, the GPR surrogate model may overly smooth the observed data if the noise variance is too high, whereas the model may overfit the noisy observations if the noise variance is too low. Both cases lead to suboptimal decisions. All three hyperparameters are autotuned in the simulation and tracked with iterations.
The next four points to sample are suggested according to the acquisition function and batch-picking method (Fig. 2(d)). We test LP, KB, and CL as batch-picking methods for both Ackley and Hartmann functions with both EI and UCB as the acquisition functions with optimized hyperparameters based on performance metrics. We find LP outperforms KB and CL for both functions independent of the choice of acquisition function. For details see Supporting Information appendix section D. For the rest of the paper, LP is used as the batch-picking method unless explicitly specified.
@@ 3.4 Metrics for Performance Evaluation During Benchmarking
[…]
@@ 4 Results and Discussion
To gain insight on the behavior of BO under conditions commonly met in experimental materials science applications, we repeated simulated benchmarks of BO under varying levels of noise for two analytical test objective functions illustrated in Fig. 1: (a) A($\bf X$) representing a needle-in-a-haystack problem that could be encountered in, e.g., molecule property optimizations, and (b) H($\bf X$) representing a landscape with a false maximum that could be encountered in, e.g., material composition optimizations. BO was compared in these scenarios with two acquisition functions (EI and UCB) with respective hyperparameters, as well as three batch-picking methods (LP, KB, and CL). These options represent typical selections that need to be made while setting up an experimental materials optimization campaign. The benchmark workflow (Fig. 2) can be used for any simulated test function that represents the researchers’ beliefs of the experimental problem landscape based on the prior expertise and experiences of the researcher on the specific application, with noise levels that are estimated to be realistic for the experimental setting. Next, the observations from the benchmarks are detailed in noise-free and noisy scenarios for the two different experiment landscapes.
4.1 Noise-free Scenario
4.1.1 Needle-in-a-Haystack Landscape, Ackley
@@
Figure 3: BO results using UCB with $\beta$ = 1 on noise-free Ackley function. Learning curve in (a) ${\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for all 99 LHS initial starts. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representation (x1 vs. x2) for the 50th percentile BO model, showing how BO iterations zero in on ${\bf X}_{\rm max}$ (cyan cross). Light blue circles are the 24 initial LHS selections and the pink to dark red points are training points (progressively darker). (d) Zoomed-in 2D heat map for variables x1 and x2 within the range [-2, 2] near ${\bf X}_{\rm max}$. (e) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
@@
Fig. 3(b) shows the $y$ value estimated by the surrogate GPR model, $\mu_{\rm D}({\bf X}^{*})$, as a function of BO iterations. At iteration 1, all $y$-value estimates are below 15-15, whereas around iteration 20 estimates have improved to be above 5-5, and at the final iteration the $y$-values are all close to 0, the maximum $A({\bf X})$ value. Fig. 3(c) shows a 3D representation of the surrogate model and scatter plot of the inputs of the 50th percentile sampling on two of the six $\bf X$} dimensions. Along the input range from -32.768 to 32.768, the objective function ranges from - 22.3 to 0. Each dot indicates an input point; blue dots indicate initial LHS points and the red dots from light to dark represents the optimization progress. As iterations progress more points are sampled close to ${\bf X}_{\rm max}$, indicating that the BO exploits the optimum region. Fig. 3(d) shows the 2D heat map with focused domain range near maximum $A({\bf X})$ and reveals that most of the points for the Ackley function are sampled around ${\bf X}_{\rm max}$.
In addition to the analysis of performance metrics, the evolution of GPR hyperparameters is provided in Fig. S3. We observe that all length scales converged to similar values as expected from the rotational symmetry around the origin of the Ackley function. It should be noted that the exploration hyperparameter $\beta$ greatly affects BO convergence.[43, 27] Table  S1 shows the different metrics, $\rm\langle IR({\bf X})\rangle$, $\rm\langle CR({\bf X})\rangle$, $\rm\langle IR(y)\rangle$, and $\rm\langle CR(y)\rangle$, all normalized to the ${\bf X},y$ ranges, for different $\beta$ values. We chose the $\beta$ that produces the smallest $\rm\langle IR(\bf X)\rangle$ which we determine as $\beta=1$, which also yields the smallest $\langle{\rm CR}({\bf X})\rangle$ and $\langle{\rm CR}(y)\rangle$.
@@
4.1.2 Landscape with Nearly Degenerate Maxima, Hartmann
@@
Fig. 5a shows ${\bf X}_{k}^{i}-{\bf X}_{\rm max}||$ for the Hartmann test function using the UCB acquisition policy. Initially, the distance to the maximum ${\bf X}^{1}_{k}-{\bf X}_{\rm max}||$ ranges from $0.2$ to $1.5$ in the unit hypercube. The distance to ${\bf X}_{\rm max}$ for the Hartmann function show relatively larger variance compared to the Ackley, which we attribute to those LHS sampling result in finding a optimum close to ${\bf X}_{\rm max,2}$ rather than ${\bf X}_{\rm max}$. In Fig. 5(a), for $\approx 75\%$ of the samplings, ${\bf X}_{k}-{\bf X}_{\rm max}||$ converges to 0 while for the remaining $\approx 25\%$, ${\bf X}_{k}-{\bf X}_{\rm max}||$ converges to ${\bf X}_{\rm max,2}-{\bf X}_{\rm max}||\approx 1.1$. In some LHS samplings ${\bf X}_{k}$ is initially near ${\bf X}_{\rm max,2}$ but still ends up at ${\bf X}_{\rm max}$. Overall, because of the second maximum, it is significantly more challenging to find the GT max in the Hartmann landscape compared to the Ackley landscape.
Figure 5: BO results using UCB with $\beta$ = 1 on noise-free Hartmann function. Learning curve in (a) ${\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for 99 LHS BO models. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representations for (x1 vs. x2) variables pair, for the 50th percentile BO model at the last iteration, showing how BO iterations zero in on ${\bf X}_{\rm max}$. Red circles are the 24 initial LHS selections and the blue to light green points are training points (blue progressively to green).(d) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
The dependence of the BO results on the UCB exploration hyperparameter $\beta$ is shown in Table  S3. Similarly to Ackley, $\beta$ = 1 produces the lowest $\rm\langle IR(y)\rangle$ when using UCB policy. Comparing Table  S1 and  S3 for $\beta$ = 1 shows that ${\rm\langle IR}({\bf X})\rangle$ and $\rm\langle IR(y)\rangle$ are $\sim$ 210 times larger and $\sim$ 2 times smaller, respectively, for the Hartmann compared to the Ackley function. Our findings are in line with the benchmarking study of Liang et al.[4]
In Fig. 5(c), the $\bf X$} evaluated at each iteration of the 50th percentile (median outcome) LHS sampling are shown superimposed on the 3D representations of the GPR model. The red dots are LHS points and the blue-to-green dots represent BO optimization progress. The projections onto the other pairs of input variables are shown Fig.  S6. The GPR model of the median sampling, hence all the percentile samplings above the median, accurately identify ${\bf X}_{\rm max}$ of the test function. The sampled points (blue to light green circles) converge on ${\bf X}_{\rm max}$ and no evaluated points appear near ${\bf X}_{\rm max,2}$ demonstrating that in this case, the BO model correctly identifies ${\bf X}_{\rm max}$. The parity plot in Fig. 5(d) compares the median BO model predictions with the GT data for all the evaluated points. RMSE value for this plot is $2.6\times 10^{-4})$. The normalized RMSE value is $7.8\times 10^{-5}$, which corresponds well to the converged square root value of the GNV hyperparameter (Supplementary Fig.  S7(h)).
Fig. S8 shows still snapshots of the BO iterations for the 50th percentile LHS sampling, projected onto the x1-x3 plane, whereas Movie (2) shows the entire iterative progress of the BO. Note that in Fig. 5(c) the 3D representation is of the GP model of the final iteration whereas in Fig. S8, the evolution of the GPR model is illustrated. Initially, the GPR model displays a very flat landscape without a peak and the $\mu_{\rm D}({\bf X}^{*})$ value is <2<2. As iterations progress, the GPR model develops a peak near the GT max ${\bf X}_{\rm max}$ (black cross). By the tenth iteration, the BO model correctly identifies the ${\bf X}_{\rm max}$. Note that there is no peak near the $2^{\rm nd}$ max ${\bf X}_{\rm max,2}$ (red star).
4.1.3 Comparison of Acquisition Function Performance
The UCB and EI acquisition policies are compared for the Ackley and the Hartmann function optimization in Table 1 for three values of the exploration hyperparameters ($\beta$ for UCB and $\xi$ for EI, respectively) each. Comprehensive studies with all exploration hyperparameter values tested can be found in Table  S1 and  S3 for UCB and Table  S2 and  S4 for EI.
Table 1: Comparison between UCB and EI acquisition policies with noise-free Ackley and Hartmann test functions. Exploration parameter values ($\beta$ for UCB and $\xi$ for EI. respectively) are also listed.
Test Function
Acquisition
Policy
$\rm\langle IR({\bf X})\rangle$
1 $\times 10^{-2}$
$\rm\langle CR({\bf X})\rangle$
$\rm\langle IR(y)\rangle$
1 $\times 10^{-2}$
$\rm\langle CR(y)\rangle$
Ackley UCB ($\beta$ = 1) 0.11 1.17 1.63 4.88
Ackley UCB ($\beta$ = 15) 12.4 10.1 33.8 21.5
Ackley UCB ($\beta$ = 30) 22.1 11.3 44.8 22.8
Ackley EI ($\xi$ = 0) 0.59 1.87 9.35 8.61
Ackley EI ($\xi$ = 0.1) 3.48 6.08 26.7 19.1
Ackley EI ($\xi$ = 10) 24.3 13.6 62.6 32.6
Hartmann UCB ($\beta$ = 1) 23.1 18.5 0.84 3.31
Hartmann UCB ($\beta$ = 15) 57.0 35.1 6.95 14.1
Hartmann UCB ($\beta$ = 30) 95.5 41.4 30.7 22.6
Hartmann EI ($\xi$ = 0) 33.0 18.9 0.91 3.10
Hartmann EI ($\xi$ = 0.1) 37.0 21.1 3.01 4.12
Hartmann EI ($\xi$ = 10) 66.0 33.9 28.9 19.3
Four BO performance evaluation metrics of averaged instantaneous and cumulative regrets (${\rm\langle IR}({\bf X})\rangle$, ${\rm\langle CR}({\bf X})\rangle$, ${\rm\langle IR}(y)\rangle$, and ${\rm\langle CR}(y)\rangle$) are compared: the lower the value, the better the performance. Regardless of the metric or objective function, UCB performs best when $\beta=1$, and EI performs best for $\xi=0$. For all except $\rm\langle CR(y)\rangle$ for Hartmann, UCB yields lower regret values than EI; in the case of Hartmann $\rm\langle CR(y)\rangle$, EI produces a slightly lower value (3.10) than UCB (3.31). Figs. S9 and  S10 clearly depict the superior performance of UCB compared to EI visually via the evolution of ${\bf X}^{*}-{\bf X}_{\rm max}||$ and $\mu_{\rm D}({\bf X}^{*})$. The different is more significant for the Ackley function (Figs. S9), with the largest difference observed in ${\rm\langle IR}({\bf X})\rangle$, where the UCB value is more than 5 times lower. Overall, we conclude that UCB significantly outperforms the EI, especially in terms of the convergence of $\mu_{\rm D}({\bf X}^{*})$, which is also in line with Liang et al.[4].
To strengthen these descriptive comparisons, we examine the results at the end of optimization (50th iteration) across 99 LHS runs comparing using the best UCB ($\beta=1$) and EI ($\xi=0$). For the Ackley function, there is no overlaps between the UCB and EI results for either ${\rm IR}({\bf X})$ and ${\rm IR}(y)$, hence no need for further statistical tests. The median ${\rm IR}({\bf X})$ obtained by UCB and EI are $0.11\times 10^{-2}$ and $0.60\times 10^{-2}$, respectively, and the median ${\rm IR}(y)$ obtained by UCB and EI are $1.4\times 10^{-2}$ and $9.1\times 10^{-2}$, respectively. For the Hartmann function, there are substantial overlaps between the results obtained by the two acquisition functions; hence, we perform non-parametric Mann–Whitney U tests on them. The test reports a statistic ($U$) that measures how separated the ranks of the two groups are, and a probability value ($p$) that quantifies the likelihood of observing such a difference by chance. For the Hartmann function, the median ${\rm IR}({\bf X})$ obtained by UCB and EI are $2.5\times 10^{-2}$ and $5.7\times 10^{-2}$, respectively, yielding a $U=2829$ and $p\approx 10^{-7}$; the median ${\rm IR}(y)$ obtained by UCB and EI are $0.22\times 10^{-2}$ and $0.44\times 10^{-2}$, respectively, yielding a $U=4019$ and $p\approx 0.016$, still smaller than the $0.05$ used for statistically not able to reject the null hypothesis. Thus, for all cases, UCB and EI yield distinct performance distributions. The statistical tests further reinforce that the acquisition strategies differ meaningfully, with UCB outperforms EI.
4.2 Noisy Scenario
4.2.1 Comparing Utility Functions in the Presence of Noise
Fig. 6a-d compares the learning curves using ${\rm Max}(y)$ (a, c) vs $\mu_{\rm D}({\bf X}^{*})$ (b, d) utility functions with 5% noise for both Ackley (a, b) and Hartmann (c, d) test functions. The slower convergence in Fig. 6 compared to Figs. 3 and 5 is due to the presence of noise in the data. The ${\rm Max}(y)$ results (left column) reveal that about 5 LHS samplings yield ${\rm Max}(y)>A({\bf X}_{\rm max})$ for Ackley, and the majority of LHS samplings produce ${\rm Max}(y)>H({\bf X}_{\rm max})$ for Hartmann. $y$ can exceed the GT maximum value because of the noise when evaluating $\bf X$} close to the maximum or second maximum. In such a case, the learning curve is tracking the outliers, i.e., data point with the largest noise value, and the problem is exacerbated at higher noise levels. In contrast, when $\mu_{\rm D}({\bf X}^{*})$ is used as the utility function in the BO, almost no LHS sampling result in $\mu_{\rm D}({\bf X}^{*})$ exceeding the GT maximum value. Furthermore, the $\mu_{\rm D}({\bf X}^{*})$ learning curve of the Ackley test function (Fig. 6b) reveals more clearly that the algorithm has not yet converged to the optimum compared to the ${\rm Max}(y)$ curve (Fig. 6a). Overall, we conclude that using $\mu_{\rm D}({\bf X}^{*})$ as the utility function for the learning curves is a more suitable measure of BO algorithm than Max(y), especially when noise is present.[43] We have tracked $\mu_{\rm D}({\bf X}^{*})$ in all other learning curves in the paper.
4.2.2 Needle-in-a-Haystack Landscape, Noisy Ackley
Fig. 7 (a), and (b), show the learning curves of the Ackley test function for noise levels of 2%, 5%, 7%, and 10%, from left to right, using EI as the acquisition policy. With 2% noise, the curves look similar compared to the no-noise case although at iteration 10, one LHS sampling does not satisfy ${\bf X}^{*}-{\bf X}_{\rm max}||<10$ whereas all satisfied this criterion for the noiseless case. At 5% noise, we observe that many samplings do not reach the GT maximum but the 50th percentile sampling still reaches a value close to ${\bf X}_{\rm max}$. This indicates that a little more than half of the samplings find the global maximum location by the end of BO despite $\mu_{\rm D}({\bf X}^{*})$ is lower than $A({\bf X}_{\rm max})$.
Figure 7: Learning curve of ${\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) for the Ackley function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0), 5% ($\xi$ = 0.05), 7% ($\xi$ = 0.1), and 10% ($\xi$ = 0.05) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (blue) and 50 iterations (from light to dark red) of BO points.
Exploration hyperparameters are investigated for both UCB and EI acquisition functions as a function of noise level, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S9 and  S10). Comparing UCB $\beta=1$ and EI $\xi=0$, we find that for noise levels above 3%, the EI acquisition performs better than UCB. The better performance using the EI acquisition function motivated us to compare BO results for different noise levels using EI in Fig. 7. Neither EI nor UCB have been initially designed to be used for noisy objectives,[50] and ${\rm\langle IR}({\bf X})\rangle$ are indeed high beyond 10% of noise, indicating most of the LHS samplings have not resulted in a converged BO. Lately, acquisition functions have been designed specifically for noisy objectives, such as Noisy-EI,[50, 51] however, these have not yet been integrated into most BO packages widely used in materials science applications. Table  S11 shows the acquisition function and hyperparameter that yields the lowest ${\rm\langle IR}({\bf X})\rangle$ for each noise level, and reveals that for 2% noise, UCB with $\beta=2$ is optimal whereas for 5% and 10%, EI with $\xi$ = 0.05 is optimal, and for 7% noise, EI with $\xi=0$ is optimal. Generally, low $\xi$ values that lead to more exploitation perform clearly better in low-noise scenarios for Ackley, whereas in high-noise scenarios, the choice of $\xi$ does not affect the result drastically. In the case of UCB, the effects of the exploration hyperparameter on the convergence are not as drastic but the low exploration hyperparameter values also perform better in low-noise scenarios. Thus, based on this benchmark, low exploration choices are a robust option for EI and UCB with Ackley-type objectives.
Fig. 7(c) shows the GPR model of the sampling with the $50^{\rm th}$ percentile ${\rm\langle IR}({\bf X})\rangle$ and Fig. 7(d) shows the corresponding parity plots. For 2% noise, the GPR model resembles the GT; for 5%, the peak is slightly degraded; for 7%, the peak is significantly degraded; and finally, for 10%, the landscape is almost flat and the Ackley peak has all but disappeared. A zoom-in 2D heat map is shown in Fig. S15, and the GNV hyperparameters are shown in Fig. S16. Supplementary Movie (5) illustrates the progress of BO learning, showcasing the transformation from the initial to the final iteration of the $50^{\rm th}$ percentile sampling with 5% noise. Overall, we observe good performance at 2% noise with performance degraded as noise increases to almost no BO effectiveness at 10% noise. It is clear from 7(d) that the GPR model misses the peak when noise is above 5%. For 10% noise, the model was only covering $y$ values below -15, significantly lower than the peak. However, for Ackley, these regions constitute almost the entire ($\langle$ 99.99%) GT hypervolume. This fact is important when we compare BO performance for Ackley vs. Hartmann in Fig. 12.
4.2.3 Landscape with Nearly Degenerate Maxima, Noisy Hartmann
Figure 8: Learning curve of ${\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) of the Hartmann function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0.005), 5% ($\xi$ = 0.1), 7% ($\xi$ = 0), and 10% ($\xi$ = 0) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (red) and 50 iterations (from blue to green) of BO points.
We investigated the optimal exploration hyperparameters for both UCB and EI acquisition functions as a function of noise, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S12 and  S13). Trends similar to the Ackley test function are found for Hartmann, albeit Hartmann test function is less sensitive to the choice. The smallest ${\rm\langle IR}({\bf X})\rangle$-yielding values are used for Fig.8 (EI with $\xi=0.005$, $0.1$, 0, 0 for 2%, 5%, 7%, and 10% noise, respectively). Table  S14 highlights the best-performing acquisition function and hyperparameter for Hartmann test function at each noise level. EI is the best acquisition function for five of the noise conditions whereas UCB is the best for six of the noise conditions with only small differences between each. This indicates that for the Hartmann landscape, UCB and EI have similar BO performance.
Fig. 9 and in Movie (6) (see supplementary) show the full iterative progress of the BO of the Hartmann function with 5% noise using the EI acquisition policy ($\xi$ = 0.1). Initially, the maximum of the GPR landscape is not located near ${\bf X}_{\rm max}$. However, at iteration 32, the optimum has been identified. In comparison, in the noise-free case (Fig.  S8), ${\bf X}_{\rm max}$ and $H({\bf X}_{\rm max})$ are identified after only 5 iterations.